# Spectrum of slip dynamics, scaling & statistical laws emerge from simplified model of fault and damage zone architecture

### v1

@@ 4.6 Apparent Diffusion
It has been observed that the migration of seismicity in some fault zones follows a specific time-space curve, similar to the one obtained from fluid diffusion [[, e.g.,]]Danr2024. In other words, the distance $r$ from the epicenter versus the time from the main shock $t$ is observed to be $r\propto D\sqrt{t}$. This may seem surprising if one attributes pore pressure diffusion in a medium with uniform hydraulic propreties as the only physical mechanism that gives rise to this scaling. In our simulations, we observe the same trend despite the absence of any underlying fluid diffusion process. This observation is illustrated in Figure 10, where two slow-slip ruptures are shown (panels a–c) along with seismic events occurring off-fault overlaid on the slip rate panels. The relative distance is calculated as the distance between the off-fault event and the detected nucleation point on the main fault, using a slip rate threshold of $10^{-9}$ m/s. The relative time is determined by the time difference between the occurrence of the off-fault event and that of the main fault. This off-fault seismicity migration is shown in panels (b-d). The values of the perceived diffusivity parameter $D$ (0.02–0.06 m2/s) fall within the observed range for the fluid diffusivity parameter [6].
@@
Our simulation thus reveals an apparent diffusive process that is driven by the slow slip process, putting a major caveat about inferring underlying fluid-driven processes just from looking at the migration of seismicity following $r\propto D\sqrt{t}$. This behavior can also emerge from the physics of interacting faults governed by rate-and-state friction along with a slow slip event occurring on the main fault.
4.7 Source Time Functions across scales
Figure 11: Moment-rate functions (MRFs) (left panels), and their spectra (right panels), across different $M_{w}$ ranges for both off-fault and main fault events. (a) Selected MRFs from slow and fast events in the off-fault region within a narrow range of $0.9\leq M_{w}\leq-0.8$. (b) Same as (a), but for $0.3\leq M_{w}\leq-0.2$. (c) MRFs from main fault events, and their spectra, with a broader range of $2.2\leq M_{w}\leq 3.8$ and $2.4\leq M_{w}\leq 4.8$, showing both slow slip and fast rupture behavior.
We present a general overview of the source time functions (STFs) generated within our fault system model, focusing on their variability in moment magnitude, duration, and spectral characteristics. The analysis includes moment-rate functions (MRFs) from both slow and fast events, considering a representative fault geometry in which off-fault fractures are oriented parallel to the main fault. We selected this geometry because it exhibits a particularly broad and diverse range of temporal behaviors, capturing both small-magnitude slow events and large slow slip events occurring on both off-faults and the main fault.
Figure 12: MRF shapes for slow and fast events after normalization by centroid time. Panels (a) and (b) correspond to slow events in the off-fault region for the same magnitude ranges as in Figures 11a and 11b. (c) shows slow events on the main fault. Panels (d) and (e) correspond to fast off-fault events with the same $M_{w}$ ranges as in (a) and (b). (f) Fast events on the main fault. Color-coded by $M_{w}$.
Figures 11a and 11b show selected MRFs from off-fault events within narrow magnitude ranges: $0.9\leq M_{w}\leq-0.8$ and $0.3\leq M_{w}\leq-0.2$, respectively. These MRFs exhibit a continuous and broad distribution of slip durations, with significant variability in their shape, ranging from abrupt moment release to smoother and more gradual evolutions. This variability highlights the influence of structural complexity and distributed deformation in controlling rupture behavior, emphasizing that seismic complexity arises not only from the main fault, as shown previously, but also from off-fault activity.
Figure 11c shows MRFs of events on the main fault, which exhibit a bimodal behavior. Due to the greater extent and roughness of the main fault, the main fault hosts the largest magnitude events, characterized by more complex STFs, including multiple acceleration and deceleration phases during the ruptures. The magnitude range considered here includes both large-magnitude, fast ruptures, partial or full, ($2.2\leq M_{w}\leq 3.8$) and long-duration slow slip events ($2.4\leq M_{w}\leq 4.8$).
The corresponding source spectra for the same events show, as expected, that the longer durations of the slow events lead to significantly lower corner frequencies than fast events across all three magnitude ranges, on both the main and off-faults. Recently, [202] analyzed the S-wave displacement spectral signature of low-frequency earthquakes in Nankai Trough, after correcting for empirically derived attenuation, and concluded that the spectra are consistent with the classical earthquake model with vastly different rupture velocities and stress drops. Our results align very well with the conclusions drawn by [202]. Beyond the corner frequency, the spectral decay of slow events appears similar to that of fast events. However, the high-frequency content is less reliably estimated, as it should be predominantly influenced by the fully dynamic response of both event types, an aspect that our quasi-dynamic model does not account for but we aim to address in future investigations. To further examine MRF characteristics across these events, we normalized the time axis of each MRF with respect to its centroid time, $T_{c}$ (Figure 12), defined following [63, 62] as:
$T_{c}=\frac{\int_{t_{1}}^{t_{2}}t\dot{M}_{0}(t)dt}{\int_{t_{1}}^{t_{2}}\dot{M}_{0}(t)dt},$ (1)
where $\dot{M}_{0}(t)$ is the moment rate, and $t_{1}$ and $t_{2}$ correspond to the onset and termination times of the MRF, defined by when the slip rate crosses a detection threshold.
One key distinction between slow and fast STFs lies in their symmetry. MRFs of slow events (Figures 12a–b-c) tend to be more asymmetric, typically exhibiting a sharp rise followed by a longer, gradual decay—indicating a positively skewed moment release. In contrast, fast earthquakes generally show more symmetric MRFs with a single, well-defined peak (Figures 12d–e-f). However, for the largest slow slip events (Figure 12c), the STFs of both fast and slow events become more irregular, and the asymmetry for some events is less distinguishable. Consistent with the findings of [119], our results suggest that fast earthquakes tend to have simpler MRFs, with fewer subevents and with a well-defined peak and gradual decay (Figures 12d–e-f). While this section focuses on a specific fault geometry, similar results were obtained for an alternative fault configuration with different off-fault orientations.
5 Sensitivity Analysis of Fault Volume Model to Fracture Orientation and Friction Parameters
Figure 13: (a) Effect of off-fault fracture orientation on the Omori law: The average Omori decay curve over a period of 30 days is calculated from the stacked curves of different fracture orientation, and the average $p$ coefficient is reported (b) Effect of off-fault fracture orientation on the b-value found from stacked catalog of fast ruptures with same off-fault fracture orientation (c) Effect of off-fault fracture orientation on the inverse Omori law (d) Effect of off-fault fracture orientation on the moment-area scaling. The $a/b$ ratio of the off-fault fractures is 0.5.
To assess the sensitivity of our fault volume model to various parameters, we conducted multiple simulations exploring different fracture orientations and frictional parameters of off-fault fractures. These simulations encompassed four main scenarios: right-lateral fractures (oriented at $\theta=150^{\circ}\pm 10^{\circ}$), left-lateral fractures (oriented at $\theta=90^{\circ}\pm 10^{\circ}$), both right and left-lateral fractures (conjugate planes at $\theta=[90^{\circ};150^{\circ}]\pm 10^{\circ}$), and fractures parallel to the major fault (oriented at $\theta=0^{\circ}\pm 10^{\circ}$). To ensure robustness, each orientation was subjected to five different statistical samplings, guaranteeing a comprehensive understanding of the system’s behavior across various geometries. Furthermore, we repeated the entire analysis, initially with an $a/b$ ratio of 0.5 for the off-fault fractures, and subsequently with a new $a/b$ ratio of 0.4, while maintaining homogeneous, rate-weakening friction conditions. In total, we tested 40 different configurations with various combinations of off-fault fracture orientations and frictional parameters. Our investigation focuses on how these parameters influence key aspects of the statistics and scaling laws for the fault volume, and by doing so what controls the slip dynamics within the fault volume.
Figure 13a illustrates the Omori law coefficient $p$ across different fracture orientations when $a/b=0.5$. Remarkably, the coefficient remains consistently around 1 across all orientations, indicating a uniform behavior regardless of these parameters. While off-fault orientation does influence the duration of the aftershock period (see the case when off-faults are aligned with the main fault), we defer this investigation to future studies as it falls outside the scope of our current research. In Figure 13b, we show the Gutenberg-Richter relationship across various fracture orientations. Remarkably, all plotted distributions conform to the Gutenberg-Richter law, displaying a similar $b$-value across orientations. Interestingly, there is no apparent influence of off-fault fracture orientation on the magnitude-frequency distribution, suggesting that this aspect does not significantly impact slip dynamics. Figure 13c shows the increasing rate of seismicity in a 3-hour window prior to the main shock for $a/b=0.5$, grouped by fracture orientation. The increasing rate of seismic events off-fault exhibits an inverse Omori law, with $p$-values close to 1.
Figure 13d demonstrates moment-rupture area scaling (for various orientations of off-fault fractures), depicting a proportional relationship between moment and rupture area raised to the power of 3/2 ($M_{0}\propto S^{3/2}$). This scaling law underscores the fault volume model’s fidelity in capturing the spatial distribution of seismic energy release across various rupture sizes. Importantly, the observation that both fast and slow ruptures exhibit almost the same scaling between moment and rupture area is noteworthy. As seismic moment is equal to stress drop multiplied by rupture area $M_{0}=\Delta\sigma S^{3/2}/C$ where $C$ is a constant and $\Delta\sigma$ is the stress drop [95], this suggests that stress drop varies at much over fewer orders of magnitudes than rupture velocity.
Our analysis across various fracture orientations and frictional parameters reveals consistent slip dynamics. Despite orientation or $a/b$ ratio variations, our fault volume model consistently reproduces the broad set of observations. We observe increased activity compared to single fault scenarios, with minimal influence from fracture orientation on recurrence time, event frequency, or Omori law coefficient. Additionally, the magnitude-frequency distribution remains consistent across orientations. These findings underscore the robustness of our model in capturing seismic processes and suggest its applicability to real-world scenarios. Furthermore, while fracture orientation may impact the intricacies of slip dynamics within the damage zone, such as the spatial distribution of foreshocks and aftershocks or the occurrence of cascading ruptures, these details are beyond the scope of our current study. Future investigations will delve into these aspects for a comprehensive understanding of fault behavior.
In Figure 14, we present a comprehensive catalog compiled from all simulations conducted, totaling around 1 million events, providing insight into the broad statistical features characterizing seismic activity within the fault volume model. Figure 14 illustrates moment-duration scaling, revealing a consistent relationship where moment is proportional to the cube of duration for both fast and slow ruptures. As discussed in Section 4.2, we once again notice the sensitivity of this scaling to the slip rate threshold, which may reflect the detection thresholds of seismograms and GPS stations in the field. From Figures 14a and 14b, we observe small events tend to have a continuum of rupture velocities; however, as events become larger, a distinct gap emerges between fast and slow ruptures. This gap mirrors observations reported by [81], suggesting events that are challenging to detect or exceedingly rare. Our numerical model’s ability to detect these events with confidence reinforces the presence of this gap, raising the possibility of a mechanical constraint where large events predominantly exhibit either fast or slow rupture velocities, rather than a continuum. This strongly suggests that both fast and slow ruptures may originate from the same mechanical model, despite their differing rupture velocities.
@@ 6 Discussion
These findings, when taken in total, highlight the fault volume framework’s capacity to capture global, fault-related scaling laws and statistics, both on the main fault and within the fractured medium. Moreover, the fault volume model accurately reproduces diverse slip dynamics, such as large slow slip events, low-frequency earthquakes (LFEs), and earthquakes, all within a unified framework. Although it is not explored here, tectonic tremors can be explained as well with our model framework as successive rupture of shear cracks as recently proposed by [204]. Our simulations also replicate the observed localization of deformation leading up to a rupture and subsequent delocalization of deformation into the fractured medium afterward, as recently reported by [97, 19]. This comprehensive representation of seismic activity underscores the importance of considering fault volume effects in seismic simulations. While several models produce one, or many, of the above-discussed features, the fault volume model stands out as the first to reproduce the entirety of observations and statistics simultaneously, demonstrating its efficacy in capturing the multifaceted nature of seismicity within fault systems. We note here this is not the first attempt in our community to expalin the divesity of slip dynamics. In the late 1970’s and early 1980’s there was a flurry of papers by [[]]Blandford1975, Nur1978, aki1979, andrews1981, hanks1981, Gusev1983 (among many others) that have proposed heterogeneities in tractions, friction and material properties play a major role in fault mechanics. Within this historical context, our work provides a natural framework to introduce traction heterogeneities through geometric complexity and reproduces a broader spectrum of slip dynamics than just earthquakes.
@@
This fault volume model achieves all of the above without the explicit incorporation of frictional heterogeneity or fluid presence within the medium, indicating its efficacy in explaining complex slip behaviors solely through geometric complexity, which is an independently measurable quantity. A notable distinction arises when comparing the behavior of a single fault with frictional heterogeneity to that of a fault volume. In the former case, frictional properties remain constant throughout seismic cycles, while with a fault volume model, traction (or stress) heterogeneities (asperities or barriers) naturally develop and disappear over various spatiotemporal scales. What may appear as a barrier, region of high-coupling for example, during one seismic cycle (few decades) may become an asperity during the next cycle, and vice-versa. This raises the question of what we mean by “barriers” in the context of geodetic observations, which typically span only a few decades. Such observations are often interpreted as evidence of frictional asperities, yet in reality they simply show that slip occurs when fault traction exceeds frictional resistance. Slip dynamics - whether creep, slow slip events (SSEs), or earthquakes - are additionally governed by the temporal evolution of stresses: they also control the nucleation length and thus determines whether the fault can accelerate to seismic slip velocities. Moreover, no matter how good observations become, it will be impossible to determine the frictional heterogeneity of a fault unless very strict model assumptions are made. Therefore, we argue that modeling the full spectrum of slip dynamics through geometric complexity offers a simpler, integrated, and comprehensive approach.
@@ 7 Conclusion
In this paper, we investigate how fault zone architecture can reproduce the observed spectrum of slip dynamics and their statistical properties. We perform quasi-dynamic simulations on randomly generated 2D fault zone models, where the fault zone consists of a main fault with self-similar roughness and a power-law size distribution of off-fault fractures spanning lengths from about 1 meter to kilometers. The off-fault fracture density decays with distance from the main fault, over a few cohesive zone length scales, following a power-law distribution. All faults are assigned spatially uniform, rate-weakening frictional properties ($a/b<1$), and the critical slip distance $d_{c}$ is scaled with fault length. We show that these ingredients are sufficient to reproduce the full spectrum of slip dynamics, from slow slip events to fast earthquakes, and their statistical properties.
@@
Our simulations reproduce many natural features, including the Gutenberg-Richter magnitude-frequency distribution, Omori and inverse-Omori behavior of seismicity rates before and after mainshocks, moment–duration and moment–area scaling relationships, and the observed localization-delocalization of seismicity around an event. We also observe that off-fault seismicity migrates in a manner similar to fluid diffusion in a homogeneous porous medium and raise a caveat emptor that square-root-time-like migration can emerge from the physics of interacting faults governed by rate-and-state friction, even in the absence of fluid diffusion.
@@
Despite its simplicity and reliance on broadly accepted assumptions, the fault volume model presented in this paper successfully reproduces many of the observed key physical and statistical features of a canonical fault zone, making it a true digital twin of a fault zone (see Figure 15). We name this digital twin, SAFAR (Système Analogue des FAilles Réelles). This breakthrough opens transformative possibilities: our synthetic catalogs can enhance machine learning algorithms for seismic hazard forecasting by providing physically consistent training data that spans the full spectrum of seismic behaviors. It also provides invaluable data to analytically model various physical mechanisms that are at the root of the observed scaling laws, statistics and slip dynamics. In addition, it can now be used to characterize the response of a fault system to various types of perturbations, such as fluid injection or earth tides, and to understand the underlying physics of earthquake triggering. Most importantly, by demonstrating that geometric complexity alone-without invoking unmeasurable frictional heterogeneities-can explain the rich fault dynamics, this work fundamentally shifts how we conceptualize and model earthquake systems, paving the way for more predictive and physically grounded seismic hazard assessment in the future.

### v2

@@ 4.7 Apparent Diffusion
It has been observed that the migration of seismicity in some fault zones follows a specific space-time evolution, similar to the one obtained from fluid diffusion [[, e.g.,]]Danr2024. In other words, the distance $r$ from the epicenter versus the time from the main shock $t$ is observed to be $r\propto D\sqrt{t}$. In our simulations, we observe the same evolution despite the absence of any underlying fluid diffusion process. This observation is illustrated in Figure 11, where two slow-slip ruptures are shown (panels a,c) along with seismic events occurring off-fault overlaid on the slip rate panels. The relative distance is calculated as the distance between the off-fault event and the detected nucleation point on the main fault, using a slip rate threshold of $10^{-9}$ m/s. The relative time is determined by the time difference between the occurrence of the off-fault event and that of the main fault. This off-fault seismicity migration is shown in panels (b,d). The values of the inferred diffusivity parameter $D$ (0.02–0.06 m2/s) fall within the observed range for the fluid diffusivity parameter [5].
@@
Our simulations reveal an apparent diffusive process driven by slow slip rather than fluids, demonstrating that seismicity migration following $r\propto\sqrt{t}$ does not necessarily imply underlying fluid-driven processes. This behavior can also emerge from the physics of interacting faults governed by rate-and-state friction along with a slow slip event occurring on the main fault.
@@ 5 Discussion
These findings, when taken in total, highlight the fault volume framework’s capacity to capture global, fault-related scaling laws and statistics, both on the main fault and within the fractured medium. Moreover, the fault volume model accurately reproduces diverse slip dynamics, slow to fast earthquakes across scales, all within a unified framework. Our simulations also replicate the observed localization of deformation leading up to a rupture and subsequent delocalization of deformation into the fractured medium afterward, as recently reported by [90, 17]. This comprehensive representation of seismic activity underscores the importance of considering fault volume effects in seismic simulations. While several existing models produce one, or many, of the above-discussed features, the fault volume model stands out as the first to reproduce the entirety of observations and statistics simultaneously, demonstrating its efficacy in capturing the multifaceted nature of seismicity within fault systems.
We note here this is not the first attempt in our community to explain the diversity of slip dynamics. In the late 1970’s and early 1980’s there was a flurry of papers by [[]]Blandford1975, Nur1978, aki1979, andrews1981, hanks1981, Gusev1983 (among many others) that have proposed heterogeneities in tractions, friction and material properties play a major role in fault mechanics. Within this historical context, our work provides a natural framework to introduce traction heterogeneities through geometric complexity and reproduces a broader spectrum of slip dynamics than just earthquakes.
@@
This fault volume model achieves all of the above without the explicit incorporation of frictional heterogeneity or fluid presence within the medium, indicating its efficacy in explaining complex slip behaviors solely through geometric complexity, which is an independently measurable quantity. A notable distinction arises when comparing the behavior of a single fault with frictional heterogeneity to that of a fault volume. In the former case, frictional properties remain constant throughout seismic cycles, while with a fault volume model, traction heterogeneities (asperities or barriers) naturally develop and disappear over various spatiotemporal scales. What may appear as a barrier, region of high-coupling for example, during one seismic cycle (few decades) may become an asperity during the next cycle, and vice-versa. This raises the question of what we mean by “barriers” or “asperities” in the context of geodetic observations, which typically span only a few decades. Such observations are often interpreted as evidence of frictional asperities, yet in reality they simply show that slip occurs when fault traction exceeds frictional resistance. Slip dynamics - whether creep, slow slip events (SSEs), broadband earthquakes or earthquakes - are additionally governed by the temporal evolution of tractions: they also control the nucleation length and thus determines whether the fault can accelerate to seismic slip velocities. Moreover, no matter how good observations become, it will be impossible to determine the frictional heterogeneity of a fault unless very strict model assumptions are made. Therefore, we argue that modeling the full spectrum of slip dynamics through geometric complexity offers a simpler, integrated, and comprehensive approach opening up a new avenue to develop a digital twin framework for fault systems.
Recent advances in computational earthquake physics have enabled the development of digital twin frameworks for fault systems. [68] developed a Bayesian inversion-based digital twin that couples physics-based models of earthquake rupture and tsunami generation with real-time seafloor observations, demonstrating feasibility of real-time forecasting for subduction zone events. Their approach combines 3D coupled acoustic-gravity wave equations with seafloor pressure sensor data to infer earthquake-induced seafloor motion and forecast tsunami propagation with quantified uncertainties. Complementing this operational focus, [1, 113] have developed computational platforms for simulating sequences of earthquakes and aseismic slip (SEAS) with explicit representation of fault zone complexity, including off-fault plasticity and secondary fractures. Their hybrid finite element-spectral boundary integral method aims toward “realizing Digital Twins for crustal volumes.” These efforts demonstrate complementary approaches to digital twins in earthquake science: real-time data assimilation for operational forecasting and high-fidelity physics-based simulation of fault zone processes across multiple spatiotemporal scales.
We adopt the latter definition of a digital twin for our physics-based fault volume modeling framework following the broader definition used in computational mechanics as “a high accuracy physics-based model” validated against observations [191]. Our framework provides a simple yet high-fidelity virtual representation reproducing the full spectrum of observed fault slip behaviors from first principles, its statistical and scaling laws, enabling hypothesis testing and scenario exploration impossible on natural fault systems. This digital twin generates synthetic data with perfectly known sources and complete observational coverage, allowing us to develop improved metrics and inverse methods for interpreting fault slip dynamics that we can then apply to real observations. Future incorporation of data assimilation will enhance predictive capabilities, but the current physics-based approach already establishes this framework as a powerful digital twin for fault slip dynamics.
While our framework demonstrates the digital twin concept for fault slip dynamics, several avenues merit further investigation. An inexhaustive list is compiled below. First, exploring a broader range of friction parameters, and laws, for both off-fault fractures and the main fault would help establish the generality of our findings across different geological settings. The power-law exponent of the off-fault fracture size distribution represents another key parameter whose systematic variation could reveal how fault network geometry controls the emergent spectrum of seismic and aseismic phenomena along with the statistical properties. While we observed that off-fault orientation influences aftershock duration—particularly when fractures align with the main fault—a comprehensive study of orientation effects remains for future work. Additionally, our current framework’s exclusive focus on shear fractures can lead to unrealistic normal traction accumulation; implementing a strategy to properly manage normal traction loading and unloading would better capture the full traction evolution during slip events. To definitively confirm broadband earthquakes, source time functions should be propagated through elastic waves to synthetic seismometers. Moving toward fully dynamic simulations would eliminate quasi-dynamic approximations and allow investigation of inertial effects, though at significantly higher computational cost. Finally, the physical origin of the observed convexity in the truncated Gutenberg-Richter distribution warrants theoretical investigation to understand whether it reflects fundamental constraints on maximum earthquake size or emerges from specific properties of our fault network geometry.
The model makes several testable predictions that can be evaluated using independent observations not directly targeted in this study, including (i) systematic localization and migration of microseismicity toward the eventual rupture plane prior to large events, (ii) threshold-dependent scaling of slow-slip duration inferred from high-resolution geodetic or seismic catalogs, (iii) characteristic spatiotemporal clustering of off-fault seismicity during slow slip, and (iv) statistical properties of source time functions across scales. High-resolution seismicity catalogs, dense nodal arrays, and improved geodetic inversions provide promising avenues to quantitatively test these predictions and potentially falsify the model. Such evaluations will help refine the fault volume framework and its applicability to real fault systems.
@@ 6 Conclusion
In this paper, we investigate how fault zone architecture can reproduce the observed spectrum of slip dynamics and their statistical properties. We perform quasi-dynamic simulations on randomly generated 2D fault zone models, where the fault zone consists of a main fault with self-similar roughness and a power-law size distribution of off-fault fractures spanning lengths from about 1 meter to kilometers. The off-fault fracture density decays with distance from the main fault, over a few cohesive zone length scales, following a power-law distribution. All faults are assigned spatially uniform, slip rate-weakening frictional properties, and the critical slip distance $d_{c}$ is scaled with fault length such that each fracture, if simulated independently, can undergo seismic cycles. We show that these ingredients are sufficient to reproduce the full spectrum of slip dynamics, from slow slip events to fast earthquakes, and their statistical properties.
@@
Our simulations reproduce all of the observed scaling and statistical properties of slip dynamics, including the Gutenberg-Richter magnitude-frequency distribution, Omori and inverse-Omori behavior of seismicity rates before and after mainshocks, moment–duration and moment–area scaling relationships, and the observed localization-delocalization of seismicity around an event. We also observe that off-fault seismicity migrates in a manner similar to fluid diffusion in a homogeneous porous medium and raise a caveat emptor that square-root-time-like migration can emerge from the physics of interacting faults governed by rate-and-state friction, even in the absence of fluid diffusion.
@@
We thus propose a geometry-driven modeling framework in which fault geometry provides a primary organizational structure for slip dynamics. Fault-zone geometry is an independently measurable and unavoidable property of natural systems across scales; in our framework, it naturally generates spatiotemporally evolving traction heterogeneities that govern nucleation, rupture velocity, and event interactions across the full spectrum of slip behaviors. This perspective helps reduce reliance on poorly constrained parameters and provides a physically grounded reference framework.
We do not intend to replace frictional or hydraulic heterogeneity models, but rather propose that geometric complexity constitutes a necessary baseline upon which these processes operate. Frictional and hydraulic heterogeneities may further modulate slip dynamics, but they act within a fault system whose stress organization is primarily shaped by geometry. This geometry-driven baseline provides a foundation within which additional complexities can be meaningfully evaluated and their relative contributions assessed. Such a framework enables the construction of digital twins of real fault systems—computational analogs that capture essential geometric and mechanical properties while generating realistic slip dynamics.
The fault volume model presented in this paper demonstrates the viability of this approach. Despite its simplicity and reliance on broadly accepted assumptions, it successfully reproduces many of the observed key physical and statistical features of fault zones (see Figure 12). We name this digital twin framework SAFAR (Système Analogue des FAilles Réelles). Our synthetic catalogs can enhance machine learning algorithms for seismic hazard forecasting by providing physically consistent training data that spans the full spectrum of seismic behaviors. The framework also provides data to analytically model various physical mechanisms that underlie the observed scaling laws, statistics, and slip dynamics. In addition, it can be used to characterize the response of a fault system to various types of perturbations, such as fluid injection or earth tides, and to understand the underlying physics of earthquake triggering. By demonstrating that geometric complexity alone can generate rich fault dynamics without requiring unmeasurable frictional heterogeneities, this work provides a novel framework for more predictive and physically grounded seismic hazard assessment.